solve ( x+3 )( x-5 )
Answer by dawny74(8)

solve ( x+3 )( x-5 )
(x + 3)x + (x + 3)-5
(x*x) + (x*3) + (x*-5) + (3*-5)
=x^2 + 3x + -5x + -15
= x^2 - 2x + - 15
Added explanation:
For this problem use the F.O.I.L. method (First, Outer, Inner, Last). This is essentially what I did, but my order on the second line was a little "out of order". So, following FOIL exactly:
(x*x) = First - first instances in each parentheses.
(x*-5) = Outer - outer instances in each parentheses
(x*3) = Inner - inner instances in each parentheses
(3*-5) = Last - last instances in each parentheses
x*x is a variable so you have two variables x will now become x^2
Then you combine your like terms 3x + -5x = -2x
the final solution could be written as it is above or
x^2 - 2x - 15
Either way the subtraction symbol signifies that the next number is negative.